# Splittable Spanning Trees and Balanced Forests in Dense Random Graphs

### v1

@@ 2.2. Algorithms for generating balanced forests
First, the simplest algorithm is splitting a full spanning tree (see Algorithm 1). This algorithm will run a single stage and either return a balanced partition sampled uniformly according to the spanning tree weight, or fail. When using it in practice, the algorithm will be run until it succeeds in finding a balanced $k$-partition. It is described in [10]. The algorithm first generates a random spanning tree. Then it generates a partition by splitting the tree into $k$ equal pieces, if that is possible. Then it accepts the partition with probability equal to the inverse of the number of spanning trees that would generate that partition. The algorithm finds $k-1$ edges in the spanning tree to split it equitably as follows. If $k-1$ such edges exist, they are unique, according to Lemma 2.1. Each edge, when removed by itself from the tree, splits the tree into two pieces whose sizes are multiples of $n$, and this property uniquely characterizes these $k-1$ edges.
@@
Next, we will examine two Markov-chain based algorithms. The up-down forest walk, described in [11], moves among forests with $N-k$ edges (see Algorithm 2). The number of times $t$ that the loop is iterated depends on the mixing time, which was shown to be $O(n^{2}\log(n))$ in the worst case [11].
@@
1:Input: a connected graph $G$ with $kn$ vertices and a forest $F$ on $G$ with $n-k$ edges.
@@
The final algorithm we describe is Reversible ReCom as described in [9] (see Algorithm 3). This is also a Markov Chain algorithm that only moves between balanced partitions, so always returns a valid balanced partition, however its mixing time is unknown and sometimes exponential.
@@
3:Select uniformly at random $2$ pieces $P_{i},P_{j}$.
@@
All of the listed Markov Chain algorithms are known to converge to a stationary distribution that selects each forest with equal probability, including both reversible ReCom and Forest ReCom[11][6][9]. However, ReCom Markov chains may not be irreducible [3], so may have multiple stationary distributions depending on the starting partition. Moreover, guarantees on the mixing times of ReCom are unknown, with some cases known to require exponential time[11].
@@ 3. Challenges when allowing slack
We now present the family of graphs that proves Theorem 1, which we now restate, highlighting potential problems with ReCom when allowing slack.
@@
There exists a family of graphs where the total spanning tree weight of partitions into $3$ pieces of sizes $\{n-1,n,n+1\}$ is exponentially higher than the total weight of partitions into $3$ pieces of size $n$. Further, there exist such graphs that are induced Hamiltonian subgraphs of the triangular lattice, and have a disconnected state space when using ReCom without slack.
@@
We will construct the family of graphs ${G_{n}}$ depicted in Figure 1. For any $n$, $G_{n}$ consists of a hexagon with $2$ edges on each side, attached to $4$ ladders. For any large enough even $n$, we can expand the ladders such that $G_{n}$ contains $3n$ vertices, with the red piece having a ladder of length $n-1$; the green piece containing $7$ vertices of the hexagon and each of its two ladders containing an additional $(n-6)/2$ vertices, for a total of $n+1$ vertices; and the blue piece containing $5$ vertices of the hexagon and an additional $n-5$ vertices on its ladder. This is always a valid almost-balanced partition into pieces of size $\{n-1,n,n+1\}$.
On the other hand, this graph can only be partitioned into pieces of size $n$ along its Hamiltonian cycle. This will follow from the fact that each ladder has fewer than $n$ vertices. Consider two adjacent vertices $u$ and $v$ on opposite sides of a ladder. If they are in the same piece, then every vertex further from the hexagon on the ladder must be in the same piece. If they are in different pieces, then every vertex on the same side as $u$ on the ladder closer to the hexagon must be in the same piece as $u$, and every vertex on the same side as $v$ on the ladder closer to the hexagon must be in the same piece as $v$. Therefore, each ladder is partitioned along the Hamiltonian cycle, and the vertices in the hexagon must be as well.
@@
This implies that when using ReCom (or its variants) without slack, no move is possible, as any piece’s position along the Hamiltonian cycle determines the positions of the other two pieces. Hence, any move splitting and recombining two of them cannot modify the partition, so ReCom fails to make any nontrivial moves. As the graph can only be partitioned along the Hamiltonian cycle, there are only $n$ balanced partitions of the graph up to relabeling.
To show that the balanced partitions of the graph have exponentially less total weight than the almost balanced partitions with slack at most 1, we need only show that there exists an almost balanced partition with exponentially higher spanning tree weight than any individual balanced partition.
Now consider one of the two smaller ladders, for example, the one colored blue in Figure 1(b). Including the two vertices on the hexagon it contains $n/2-1$ vertices. In any balanced partition, one piece must contain at least half of the vertices this ladder, let us say $n/4+k$ vertices, of which the last $k$ vertices form a ladder. The same piece must also contain at most $3n/4-k$ vertices on the blue or red ladder (or both), of which at most $n/2-k$ form a ladder. Therefore, at least $n/2$ vertices in this piece are not part of a ladder that is contained in this piece. By Lemma 3.1, this piece must have spanning tree weight that is exponentially smaller than any piece of the almost-balanced partition in Figure 1(a). Since this almost balanced partition contains all but $O(1)$ vertices in every ladder, it has at most a constant fraction smaller spanning tree weight from each ladder than any balanced partition, and exponentially more in at least one ladder, proving that in all it has exponentially higher weight than any balanced partition. $\Box$
@@ 4. Splittability of Dense Graphs
As Section 3 demonstrates, even the smallest amount of slack can cause balanced partitions to be exponentially supressed at stationarity. This potentially poses a problem for any algorithms sampling spanning trees until a splittable one is found. In this section we focus on Algorithms 1 and 2, that are efficient only when balanced partitions (or balanced forests) are not too unlikely compared to the weight when they balance is not required.
@@ 4.1. Splittability of the complete graph
Figure 2. Total spanning tree weights of all Partitions of size $i,N-i$ on the Complete Graph with $N=8$ (a) and $N=16$ (b).
(a)                                                                   (b)
@@
In the next subsection we will extend this result to dense random graphs. We state and prove this result first because we will use it in the proof of the stronger result.
Recall Theorem 2, which says:
@@
The number of spanning trees in $K_{N}$ is $N^{N-2}$. Each splittable spanning tree can be split uniquely into balanced pieces (Lemma 2.1); therefore, each one is uniquely defined by the partition it creates, the spanning trees on the $K_{n}$s in those partitions, and the $k-1$ edges between partitions. There are ${(kn)!}/{(n!)^{k}k!}$ such partitions. The edges between the partitions form a spanning tree on $k-1$ vertices, with each edge having $n^{2}$ possibilities. Then, by Stirling’s formula [1], the probability a tree is splittable is
$\frac{(n^{n-2})^{k}\frac{(n^{2})^{k-1}k^{k-2}}{k!}\frac{(kn)!}{(n!)^{k}}}{{(kn)}^{kn-2}}=$ $(1+O(\frac{1}{n}))^{k}(1+O(\frac{1}{k}))(1+O(\frac{1}{kn}))n^{-\frac{k-1}{2}}(\frac{e}{\sqrt{2\pi}})^{k}$
@@
If we set $C_{k}=(1+o(k))(\frac{e}{\sqrt{2\pi}})^{k}$, the theorem follows. ∎
We note that our bound for the probability that a random spanning tree is $k$-splittable on the compete graph is $\Omega(1/n^{(k-1)/2})$, better than the best known bound of $\Omega(n^{2-2k})$ on the grid. However, Algorithm 1, which generates a $k$-partition from a spanning tree, uses rejection sampling twice; once when determining if a spanning tree is $k$-splittable, and once to account for the fact that each $k$-balanced partition may be generated by many different spanning trees. This second step of rejection sampling is maximally inefficient in the complete graph, with an acceptance probability of $1/(k^{k-2}n^{2k-2})$.
@@ 4.2. Dense Random Graphs
While each labeled tree is equally likely to appear in a random graph, it is not true that they are equally likely to be selected as a random spanning tree of a random graph. The probability that any given tree is selected as a random spanning tree is proportional to $\mathbf{E}[1/\tau(G)\,|\,T\in G]$, which varies for different trees. Trees that occur in graphs that contain many other trees are less likely to be selected.
For example, a star with three leaves is more likely than a path of three edges to be selected as a spanning tree of a random graph on four vertices. We can see that adding $1$ edge to a path of $4$ vertices can create a graph with $3$ or $4$ spanning trees, while adding $1$ to a star graph always creates a graph with $3$ spanning trees. Since adding $2$ or more edges always creates $K_{4}$ or $K_{4}$ minus one edge, we see that random graphs containing a given star graph have fewer spanning trees than those containing given path graph, so stars appear more often in random spanning tres of random graphs on $4$ vertices than they do in random labeled trees on $4$ vertices. Based on numerical results, this same trend persists into star and path graphs on $n$ vertices.
Recall Theorem 3, which says:
@@
Let $N=nk$. Let $G$ be a random connected graph of $m=p\binom{N}{2}$ edges on $N$ nodes. Let $T$ be a random spanning tree of $G$. If $m=\Omega(\binom{N}{2}/\log n)$, then the probability that $T$ is $k$-splittable is at least $C_{k}n^{-\frac{k}{2}-O(1)}$, where $C_{k}$ depends on $k$. More precisely, the probability is at least $C_{k}n^{\frac{1-k}{2}}pe^{-(2+o(1))/p}$, an increasing function of $p$.
@@
Let $\tau_{k}(G)$ be the number of spanning trees of $G$ that are $k$-splittable. Let $\tau_{max}(m,N)$ be the maximum number of spanning trees in a graph of $m$ edges on $N$ vertices. For a fixed $G$ the probability that $T$ is $k$-splittable is
@@
For random $G$ it is the expected value of this ratio,
$\mathbf{E}\left[\frac{\tau_{k}(G)}{\tau(G)}\right]$ $\geq\frac{1}{\tau_{max}(m,N)}\mathbf{E}[\tau_{k}(G)]$
$=\frac{1}{\tau_{max}(m,N)}\sum_{T~{}k\mathrm{-splittable}}\mathbf{P}[T\in G]$
@@
Grimmett [14] has shown that $\tau_{max}(m,N)\leq\frac{1}{N}(\frac{2m}{N-1})^{N-1}=(pN)^{N-2}$; therefore, the probability that $T$ is $k$-splittable is at least
$\frac{C_{k}}{(pN)^{N-2}}N^{N-2}n^{\frac{1-k}{2}}\mathbf{P}[T\in G]=\frac{C_{k}}{p^{N-2}}~{}n^{\frac{1-k}{2}}~{}\mathbf{P}[T\in G],$
@@
It will now suffice to show that $\mathbf{P}[T\in G]\geq p^{N-1}e^{-(2+o(1))/p}$, because $pe^{-(2+o(1))/p}=n^{-O(1)}$. If we remove the condition that $G$ be connected, $\mathbf{P}[T\in G]$ only gets smaller, because disconnected graphs have no spanning trees. The probability that any combination of $N-1$ edges is in $G$ is
$\prod_{i=0}^{N-2}\frac{m-i}{\binom{N}{2}}$ $\geq\left(\frac{m-N+2}{\binom{N}{2}}\right)^{N-1}$
@@
Because the probability of being connected increases with $p$ we can prove the lemma for an arbitrary random graph $G$ in the $G(N,p)$ model. We will show that for each $j\leq pN$, $G$ is at least half as likely to have $pN+j$ edges as $pN-j$ edges. Letting $m=pN$,
$\frac{\mathbf{P}[|E(G)|=m+j]}{\mathbf{P}[|E(G)|=m-j]}$ $=\frac{\binom{N}{m+j}(\frac{m}{N})^{m+j}(1-\frac{m}{N})^{N-m-j}}{\binom{N}{m-j}(\frac{m}{N})^{m-j}(1-\frac{m}{N})^{N-m+j}}$
$=\frac{(N-m+j)\cdots(N-m-j+1)}{(N-m)^{2j}}\frac{m^{2j}}{(m+j)\cdots(m-j+1)}$
$=\frac{1+\frac{j}{N-m}}{1+\frac{j}{m}}\prod^{j-1}_{i=1}\frac{1-(\frac{i}{N-m})^{2}}{1-(\frac{i}{m})^{2}}$
@@
Let $N=nk$. Let $G$ be a random graph in the $G(N,p)$ model conditioned on being connected. Let $T$ be a random spanning tree of $G$. If $p=\Omega(1/\log n)$, then the probability that $T$ is $k$-splittable is at least $C_{k}n^{-\frac{k}{2}-O(1)}$, where $C_{k}$ depends only on $k$. More precisely, the probability is at least $C_{k}n^{\frac{1-k}{2}}pe^{-(2+o(1))/p}$.
@@
For any fixed $m\geq p^{\prime}\binom{N}{2}$, if we condition on $G$ having $m$ edges, then both $G(N,m)$ and $G(N,p)$ have the uniform distribution over connected graphs. Applying Theorem 4 for each such $m$, the probability that $T$ is $k$-splittable is at least
$C_{k}n^{\frac{1-k}{2}}pe^{-(2+o(1))/p}=C_{k}n^{-\frac{k}{2}-O(1)},$
where we have absorbed $1/3$ into $C_{k}$. We note that the bound in Theorem 4 is an increasing function of $p$, so we may choose the constant $O(1)$ in the exponent uniformly over $m$. The theorem follows. ∎

### v2

@@ 2.2. Algorithms for generating balanced forests
First, the simplest algorithm is splitting a full spanning tree (see Algorithm 1). This algorithm will run a single stage and either return a balanced partition sampled uniformly according to the spanning tree weight, or fail. When using it in practice, the algorithm will be run until it succeeds in finding a balanced $k$-partition. It is described in [cannon2024grid]. The algorithm first generates a random spanning tree. Then it generates a partition by splitting the tree into $k$ equal pieces, if that is possible. Then it accepts the partition with probability equal to the inverse of the number of spanning trees that would generate that partition. The algorithm finds k1k-1 edges in the spanning tree to split it equitably as follows. If k1k-1 such edges exist, they are unique, according to Lemma 2.1. Each edge, when removed by itself from the tree, splits the tree into two pieces whose sizes are multiples of $n$, and this property uniquely characterizes these k1k-1 edges.
@@
Next, we will examine two Markov-chain based algorithms. The up-down forest walk, described in [charikar2022complexity], moves among forests with NkN-k edges (see Algorithm 2). The number of times $t$ that the loop is iterated depends on the mixing time, which was shown to be $O(n^{2}\log(n))$ in the worst case [charikar2022complexity].
@@
1:Input: a connected graph $G$ with $kn$ vertices and a forest $F$ on $G$ with nkn-k edges.
@@
The final algorithm we describe is Reversible ReCom as described in [revrecom] (see Algorithm 3). This is also a Markov Chain algorithm that only moves between balanced partitions, so always returns a valid balanced partition, however its mixing time is unknown and sometimes exponential.
@@
3:Select uniformly at random 22 pieces $P_{i},P_{j}$.
@@
All of the listed Markov Chain algorithms are known to converge to a stationary distribution that selects each forest with equal probability, including both reversible ReCom and Forest ReCom[charikar2022complexity][DukeReCom1][revrecom]. However, ReCom Markov chains may not be irreducible [akitaya2022reconfiguration_recombination], so may have multiple stationary distributions depending on the starting partition. Moreover, guarantees on the mixing times of ReCom are unknown, with some cases known to require exponential time[charikar2022complexity].
@@ 3. Challenges when allowing slack
We now present the family of graphs for which the probability of a random spanning tree being splittable on this family is exponentially small, and which highlight potential problems with ReCom when allowing slack.
Definition 3.1.
A tree is $k$-splittable with slack $j$ if there exist k1k-1 edges that, when removed, create a graph whose $k$ connected components have between njn-j and $n+j$ vertices, inclusive, for some $n$. Let $\tau^{j}(G)$ be the number of splittable spanning trees of $G$ with slack $j$.
@@
There exists an infinite family $\{G_{n}:n\geq 6,n~\mathrm{even}\}$ of graphs on $3n$ vertices such that each graph $G_{n}$
  1. (1)
    is a Hamiltonian induced subgraph of the triangular lattice
  2. (2)
    contains exactly $n$ partitions into three pieces of size $n$, namely the partitions of the $3n$ Hamiltonian paths
  3. (3)
    is such that $\tau^{0}(G_{n})/\tau^{1}(G_{n})$ is exponentially small.
Corollary 3.1.
There exists an infinite family of graphs on $3n$ vertices, $n\geq 6$, for which ReCom without slack has a disconnected state space and ReCom with slack cannot sample partitions of equal weight in polynomial time.
@@
We will construct the family of graphs $G_{n}$} depicted in Figure 1. For any $n$, $G_{n}$ consists of a hexagon with $2$ edges on each side, attached to $4$ ladders. For any even $n$, $n\geq 6$, we can expand the ladders such that $G_{n}$ contains $3n$ vertices, with the red piece having a ladder of length n1n-1; the green piece containing $7$ vertices of the hexagon and each of its two ladders containing an additional (n6)/2(n-6)/2 vertices, for a total of $n+1$ vertices; and the blue piece containing $5$ vertices of the hexagon and an additional n5n-5 vertices on its ladder. This is always a valid almost-balanced partition into pieces of size $\{n-1,n,n+1\}$. This graph contains a Hamiltonian cycle that traverses the hexagon clockwise, taking a clockwise detour out and back along each ladder.
To show property (2) of the theorem, we claim that the only way to split this graph into three connected pieces of size $n$ is to split the Hamiltonian cycle into three paths of size $n$. This will follow from the fact that each ladder has fewer than $n$ vertices. Consider any way to partition $G_{n}$ into three connected pieces, $A,B,C,$ of size $n$. If one of the pieces, $A$, is not a path along the Hamiltonian cycle, then it consists of at least two disconnected paths along the Hamiltonian cycle. Traversing the Hamiltonian cycle clockwise, between the last vertex of one of these paths and the first vertex of the next path, there is a nonempty set of vertices $S$ in $B\cup C$. $S$ cannot contain a vertex of the hexagon, because in that case $A$ would be disconnected. But $S$ also cannot be a subset of one of the ladders, because in that case $|S|<n$ and $S$ would be disconnected from the rest of $B\cup C$. This is a contradiction.
We now show property (3) of the theorem. Given a partition, say a ladder is “monochromatic” if it is contained in one piece of the partition. Consider the almost balanced partition shown in Figure 1(a). There are four maximal monochromatic ladders that contain 3n63n-6 vertices in total.
Now consider any balanced partition $P$ of $G_{n}$ into induced subgraphs $P_{1},P_{2},P_{3}$. Consider the subgraph $L_{i}$ of $P_{i}$ induced by the maximal ladders in $P_{i}$. By the discussion above, $P_{i}$ consists of components of $L_{i}$, connected in a chain by paths. Therefore, $\tau(P_{i})$ = $\tau(L_{i})$. By Lemma 3.1 it will be enough to show that the monochromatic ladders of $P$ contain $3n-\Omega(n)$ vertices total.
There are four ladders and only three pieces, so either there are fewer than four monochromatic ladders, in which case we are done, or one piece $P_{i}$ of the partition contains two ladders. Those two ladders must be adjacent in the clockwise traversal of the Hamiltonian cycle. Any pair of adjacent ladders in $G_{n}$ contain $n+\Omega(n)$ total vertices, of which $P_{i}$ must be missing $\Omega(n)$. ∎
Proof Of Corollary.
@@
This implies that when using ReCom (or its variants) without slack, no move is possible, as any piece’s position along the Hamiltonian cycle determines the positions of the other two pieces. Hence, any move splitting and recombining two of them cannot modify the partition, so ReCom fails to make any nontrivial moves. As the graph can only be partitioned along the Hamiltonian cycle, there are only $n$ balanced partitions of the graph up to relabeling. We also see that when allowing any slack, we must allow the partitions with the smallest amount of slack: those with piece sizes n1,n,n+1n-1,n,n+1. Then any sample made from ReCom with slack is exponentially more likely to unbalanced than balanced, the example high weight unbalanced partition described in theorem 3 is clearly reachable as a ReCom move from a balanced partition along the Hamiltonian cycle. ∎
@@ 4. Splittability of Dense Graphs
Figure 2. Total spanning tree weights of all Partitions of size i,Nii,N-i on the Complete Graph with $N=8$ (a) and $N=16$ (b).
(a)                        (b)
As Section 3 demonstrates, even the smallest amount of slack can cause balanced partitions to be exponentially suppressed at stationarity. This potentially poses a problem for any algorithms sampling spanning trees until a splittable one is found. In this section we focus on Algorithms 1 and 2, that are efficient only when balanced partitions (or balanced forests) are not too unlikely compared to the weight when they balance is not required.
@@ 4.1. Splittability of the complete graph
Balanced forests tend not to be the mode of a histogram showing the probabilities of splits of various sizes. On the complete graph, balanced splits contain the fewest forests out of all possible partition sizes, as shown in Figure 2. This means that standard techniques showing that this distribution is log-concave or unimodal will certainly fail and the calculations need to be more precise.
In the next subsection we will extend this result to dense random graphs. We first prove our result on the complete graph because we will use it in the proof of the stronger (more general) result that follows. Recall Theorem 1, given as stated in the introduction:
@@
By Cayley’s Theorem[Cayley] the number of spanning trees in $K_{N}$ is $N^{N-2}$. Each splittable spanning tree can be split uniquely into balanced pieces (Lemma 2.1); therefore, each one is uniquely defined by the partition it creates, the spanning trees on the $K_{n}$s in those partitions, and the k1k-1 edges between partitions. The number of unlabeled $k$-partitions of $N$ vertices is
$\frac{\binom{N}{n,n,..,n}}{k!}\ =\ \frac{(kn)!}{(n!)^{k}k!}.$
The edges between the pieces in the partition can be considered a spanning tree on a separate graph whose nodes are the $k$ pieces of the partition. By another application of Cayley’s Theorem, there are $k^{k-2}$ such spanning trees. There are $n^{2}$ ways to choose each of the k1k-1 edges, because each edge connects two components of size $n$. So, given the partition, there are $k^{k-2}(n^{2})^{k-1}$ ways to choose the edges between the pieces. Again by Cayley’s Theorem, the number of ways to choose spanning trees of $k$ copies of $K_{n}$ is $(n^{n-2})^{k}$. Putting these together, the number of $k$-splittable spanning trees $\tau_{k}(K_{n})$ of $K_{N}$ is
$\tau_{k}(K_{n})\ =\ (n^{n-2})^{k}\frac{(n^{2})^{k-1}k^{k-2}}{k!}\frac{(kn)!}{(n!)^{k}}.$
Dividing by the number of spanning trees in $K_{N}$ and applying Stirling’s formula  [abramowitz] to each of the factorials in this expression, the probability a spanning tree is splittable is
$\frac{\tau_{k}(K_{n})}{{(kn)}^{kn-2}}\$$=\ \left((1+O(\frac{1}{n}))\sqrt{2\pi n}(\frac{e}{n})^{n})^{k}~(1+O(\frac{1}{k})\right)$
$\hskip 50.58878pt\times\left(\sqrt{2\pi k}(\frac{e}{k})^{k}(1+O(\frac{1}{kn}))(\sqrt{2\pi kn}(\frac{kn}{e})^{kn}\right)\left(\frac{(n^{n-2})^{k}(n^{2})^{k-1}k^{k-2}}{{(kn)}^{kn-2}}\right)$
@@
Setting $C_{k}=(1+o(k))(\frac{e}{\sqrt{2\pi}})^{k}$, the theorem follows. ∎
We note that our bound for the probability that a random spanning tree is $k$-splittable on the compete graph is $\Omega(1/n^{(k-1)/2})$, better than the best known bound of $\Omega(n^{2-2k})$ on the grid. However, Algorithm 1, which generates a $k$-partition from a spanning tree, uses rejection sampling twice; once when determining if a spanning tree is $k$-splittable, and once to account for the fact that each $k$-balanced partition may be generated by many different spanning trees. This second step of rejection sampling is inefficient in the complete graph, with an acceptance probability of $1/(k^{k-2}n^{2k-2})$.
@@ 4.2. Dense random graphs
While each spanning tree (with the nodes labeled) is equally likely to appear in a random graph, it is not true that any two spanning trees are equally likely to be selected as a random spanning tree of a random graph. The probability that any given tree is selected as a random spanning tree is proportional to $\mathbf{E}[1/\tau(G)\,|\,T\in G]$, which varies for different trees. Trees that occur in graphs that contain many other trees are less likely to be selected.
For example, it is straightforward to show that if we choose a random graph on four vertices, then choose a random spanning tree of the graph, a star with three leaves is more likely than a path of three edges. In short, this is because random graphs containing a given star graph have fewer spanning trees than those containing given path graph. Based on numerical results, this same trend persists into star and path graphs on $n$ vertices.
Recall Theorem 2, which states:
@@
Let $N=nk$. Let $G$ be a random connected graph of $m=p\binom{N}{2}$ edges on $N$ nodes. Let $T_{0}$ be a random spanning tree of $G$. If $m=\Omega(\binom{N}{2}/\log n)$, then the probability that $T_{0}$ is $k$-splittable is at least $C_{k}n^{-\frac{k}{2}-O(1)}$, where $C_{k}$ depends on $k$. More precisely, the probability is at least $C_{k}n^{\frac{1-k}{2}}e^{-(2+o(1))/p}$, an increasing function of $p$.
@@
Let $\tau_{k}(G)$ be the number of spanning trees of $G$ that are $k$-splittable. Let $\tau_{max}(m,N)$ be the maximum number of spanning trees in a graph of $m$ edges on $N$ vertices. For a fixed $G$ the probability that $T_{0}$ is $k$-splittable is
@@
For random connected $G$, the probability that a random spanning tree is connected is the expected value of this ratio,
$\mathbf{E}\left[\frac{\tau_{k}(G)}{\tau(G)}\,|\,G\text{ connected}\right]\ \geq\ \frac{1}{\tau_{max}(m,N)}\mathbf{E}[\tau_{k}(G)\,|\,G\text{ connected}]$
We can write the right-hand side of this inequality as a sum over spanning trees $T$ of $K_{N}$, and note that when removing the condition that $G$ be connected, $\mathbf{P}[T\in G]$ only gets smaller, because disconnected graphs have no spanning trees. Then
$\mathbf{E}\left[\frac{\tau_{k}(G)}{\tau(G)}\,|\,G\text{ connected}\right]$ $\ \geq\ \frac{1}{\tau_{max}(m,N)}\sum_{T~k\mathrm{-splittable}}\mathbf{P}[T\in G\,|\,G\text{ connected}]$
$\ \geq\ \frac{1}{\tau_{max}(m,N)}\sum_{T~k\mathrm{-splittable}}\mathbf{P}[T\in G]$
@@
Grimmett [grimmett_upper_1976] has shown that $\tau_{max}(m,N)\leq\frac{1}{N}(\frac{2m}{N-1})^{N-1}=p^{N-1}N^{N-2}$; therefore, the probability that our tree $T_{0}$ is $k$-splittable is at least
$\frac{C_{k}}{p^{N-1}N^{N-2}}N^{N-2}n^{\frac{1-k}{2}}\mathbf{P}[T\in G]\ =\ \frac{C_{k}}{p^{N-1}}~n^{\frac{1-k}{2}}~\mathbf{P}[T\in G],$
@@
It will now suffice to show that $\mathbf{P}[T\in G]\geq p^{N-1}e^{-(2+o(1))/p}$, because $pe^{-(2+o(1))/p}=n^{-O(1)}$. The probability that any set $\{e_{1},...,e_{N-1}\}$ of N1N-1 edges is in $G$ is
$\prod_{i=1}^{N-1}\mathbf{P}[e_{i}\in G\,|\,e_{j}\in G\,\,\forall j<i]\$$=\ \prod_{i=0}^{N-2}\frac{m-i}{\binom{N}{2}-i}$
$\ \geq\ \left(\frac{m-N+2}{\binom{N}{2}}\right)^{N-1}$
@@
Because the probability of being connected increases with $p$ we can prove the lemma for an arbitrary random graph $G$ in the $G(N,p)$ model. We will show that for each $j\leq p\binom{N}{2}$, $G$ is at least half as likely to have $p\binom{N}{2}+j$ edges as $p\binom{N}{2}-j$ edges. Letting $M\ =\ \binom{N}{2}$ and $m=pM$,
$\frac{\mathbf{P}[|E(G)|=m+j]}{\mathbf{P}[|E(G)|=m-j]}$ $\ =\ \frac{\binom{M}{m+j}(\frac{m}{M})^{m+j}(1-\frac{m}{M})^{M-m-j}}{\binom{M}{m-j}(\frac{m}{M})^{m-j}(1-\frac{m}{M})^{M-m+j}}$
$\ =\ \frac{(M-m+j)\cdots(M-m-j+1)}{(M-m)^{2j}}\frac{m^{2j}}{(m+j)\cdots(m-j+1)}$
$\ =\ \frac{1+\frac{j}{M-m}}{1+\frac{j}{m}}\prod^{j-1}_{i=1}\frac{1-(\frac{i}{M-m})^{2}}{1-(\frac{i}{m})^{2}}$
@@
Let $N=nk$. Let $G$ be a random graph in the $G(N,p)$ model conditioned on being connected. Let $T$ be a random spanning tree of $G$. If $p=\Omega(1/\log n)$, then the probability that $T$ is $k$-splittable is at least $C_{k}n^{-\frac{k}{2}-O(1)}$, where $C_{k}$ depends only on $k$. More precisely, the probability is at least $C_{k}n^{\frac{1-k}{2}}e^{-(2+o(1))/p}$.
@@
For any fixed $m\geq p^{\prime}\binom{N}{2}$, if we condition on $G$ having $m$ edges, then both $G(N,m)$ and $G(N,p)$ have the uniform distribution over connected graphs. We now apply Theorem 4 for each such $m$. Recalling that the bound in Theorem 4 is an increasing function of $m/\binom{N}{2}$, the probability that $T$ is $k$-splittable is at least
$C_{k}n^{\frac{1-k}{2}}e^{-(2+o(1))/p^{\prime}}=C_{k}n^{-\frac{k}{2}-O(1)},$
where we have absorbed the constants $1/3$ and $p^{\prime}/p$ into $C_{k}$. The theorem follows. ∎